# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 3.1. GRT and its properties
Let $\Phi_{1}(x,y):\mathcal{U}\times\mathcal{V}\to\mathbb{R}^{2}$ be a defining function for a GRT $\mathcal{R}$. An open set $\mathcal{U}\subset\mathbb{R}^{2}$ is the image domain. We assume $y=(\alpha,p)$, and $\mathcal{V}=I_{\alpha}\times I_{p}$ is the data domain.
@@
    $I_{\alpha}\subset\mathbb{R}$ is a compact interval with the endpoints identified (so it is topologically equivalent to a circle),
@@
    $I_{p}\subset\mathbb{R}$ is an open interval (or all of $\mathbb{R}$), and
@@
    $\Phi_{1}(x,y)=p-\Phi(x,\alpha)$, $\Phi\in C^{\infty}(\mathcal{U}\times I_{\alpha})$.
Both $\mathcal{U}$ and $\mathcal{V}$ are endowed with the usual Euclidean metric. By a partition of unity we will always identify subsets of $\mathcal{V}$ with subsets of $\mathbb{R}^{2}$. The transform $\mathcal{R}$ integrates over smooth curves $\mathcal{S}_{(\alpha,p)}:=\{x\in\mathcal{U}:p=\Phi(x,\alpha)\}$:
@@
where $W\in C^{\infty}(\mathcal{U}\times\mathcal{V})$.
Here and everywhere below we use the convention that whenever $y$, $\alpha$, and $p$ appear in the same equation or sentence, then $y=(\alpha,p)$. We assume $f$ is compactly supported, $\text{supp}f\subset\mathcal{U}$, and $f$ is sufficiently smooth, so $g(y):=\mathcal{R}f(y)$ is a continuous function. Assumption 3.1 implies that $\text{supp}\,(\mathcal{R}f)$ is compact if $\text{supp}\,f$ is compact.
@@
(3.2) $\Sigma:=\{(y,x,\nu)\in\mathcal{V}\times\mathcal{U}\times\mathbb{R}:\ p-\Phi(x,\alpha)=0\}.$
Since $\partial_{\nu}(\nu\Phi_{1})=\Phi_{1}$ and $\text{d}_{(x,y,\nu)}\Phi_{1}\not=0$, the phase $\nu\Phi_{1}(x,\alpha)$ is clean and nondegenerate [21, Definition 21.2.15]. The phase $\nu\Phi_{1}(x,\alpha)$ determines a homogeneous canonical relation from $T^{*}\mathcal{U}$ to $T^{*}\mathcal{V}$ with the graph
(3.3) $\begin{split}\mathcal{C}:=&\{(y,\nu(-\Phi_{\alpha}^{\prime}(x,\alpha),1));(x,\nu\Phi_{x}^{\prime}(x,y)):(y,x,\nu)\in\Sigma\}.\end{split}$
Also, $\mathcal{R}\in I^{-1/2}(\mathcal{V}\times\mathcal{U},\mathcal{C})$ is an FIO of order $-1/2$ from $\mathcal{U}$ to $\mathcal{V}$ associated with $\mathcal{C}$ [53, Section 5.VIII]. Hence $\mathcal{R}:H_{c}^{s}(\mathcal{U})\to H_{loc}^{s+(1/2)}(\mathcal{V})$ is continuous for any $s\in\mathbb{R}$ [22, Corollary 25.3.2].
Next we compute the following determinant [53, equations (6.3) and (6.4), Chapter VIII]:
(3.4) $\begin{split}\Delta_{\Phi}(x,y,\nu):=&\text{det}\begin{pmatrix}[\nu\Phi_{1}(x,\alpha)]_{xy}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{x\nu}^{\prime\prime}\\ [\nu\Phi_{1}(x,\alpha)]_{\nu y}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{\nu\nu}^{\prime\prime}\end{pmatrix}\\ =&\text{det}\begin{pmatrix}-\nu\Phi_{x\alpha}^{\prime\prime}&(0,0)^{T}&-\Phi_{x}^{\prime}\\ -\Phi_{\alpha}^{\prime}&1&0\end{pmatrix}\\ =&\nu\text{det}\begin{pmatrix}\text{d}_{x}\Phi&\text{d}_{x}(\partial_{\alpha}\Phi)\end{pmatrix}.\end{split}$
Here $\Phi_{x}^{\prime}=(\partial_{x_{1}}\Phi,\partial_{x_{2}}\Phi)^{T}$, and all the other derivatives are defined similarly. The derivatives on the right are computed at $(x,\alpha,p=\Phi(x,\alpha))$. As is easily seen,
(3.5) $\Delta_{\Phi}(x,y,\nu)=\nu\Delta_{\Phi}(x,\alpha).$
@@
    For each $(x,\xi)\in T^{*}\mathcal{U}$, there exist $\alpha\in I_{\alpha}$ and $\nu\in\mathbb{R}$ such that $\xi=\nu\text{d}_{x}\Phi(x,\alpha)$.
@@
    $W(x,y)>c$ on $\mathcal{U}\times\mathcal{V}$.
@@
    For any bounded domain $\mathcal{U}_{b}\subset\mathcal{U}$ there exists $c=c(\mathcal{U}_{b})>0$ such that
    (3.6) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})}\text{for any }f\in H_{c}^{-1/2}(\mathcal{U}_{b}).$
Assumption 3.2(1) implies that $\Delta_{\Phi}(x,y,\nu)\not=0$ on $\mathcal{U}\times\mathcal{V}\times(\mathbb{R}\setminus 0)$. In turn, this implies that $\mathcal{C}$ is the graph of a local diffeomorphism [53, Section VI.4]. Assumption 3.2(2) implies that the natural projection $\mathcal{C}\to T^{*}\mathcal{V}$ is an embedding (injective immersion). Assumption 3.2(3) asserts that any singularity of $f$ is visible from the data. Assumption 3.2(5) implies that $\mathcal{R}$ is boundedly invertible in the scale of Sobolev spaces.
@@ 3.2. The function to be reconstructed
(3.8) $\begin{split}&H\in C^{\infty}(\mathcal{U}),\quad\text{d}H\not=0\text{ on }\mathcal{U};\\ &\tilde{f}(x,\lambda)=-i\tilde{f}_{0}(x)\lambda^{-1}+\tilde{R}(x,\lambda),\forall x\in\mathcal{U},|\lambda|\geq 1;\\ &\tilde{f}(x,\lambda)\equiv 0\ \forall x\in\mathcal{U}\setminus\mathcal{U}_{b}\text{ for some bounded domain }\mathcal{U}_{b}\subset\mathcal{U};\\ &\tilde{f}\in S^{-1}(\mathcal{U}\times\mathbb{R}),\ \tilde{f}_{0}\in C_{0}^{\infty}(\mathcal{U}_{b}),\ \tilde{R}\in S^{-2}(\mathcal{U}\times\mathbb{R}),\end{split}$
@@
Thus $\mathcal{S}$ is a smooth curve (submanifold) with possibly several connected components.
Equations (3.7) and (3.8) imply that $f\in I^{-1}(\mathcal{U},\mathcal{S})$ is a conormal distribution [21, Section 18.2] and
(3.10) $\begin{split}&f(x+\Delta x)\sim\tilde{f}_{0}(x)(\text{sgn}h/2),\ h:=\text{d}H(x)\Delta x,|\Delta x|\to 0,\ \forall x\in\mathcal{S}.\end{split}$
Thus, $\tilde{f}_{0}(x)$ is the value of the jump of $f$ across $\mathcal{S}$ at $x\in\mathcal{S}$.
@@
(3.11) $\varkappa_{\mathcal{S}}(x)=-\frac{(H_{xx}^{\prime\prime}(x)e,e)}{|H_{x}^{\prime}(x)|},\ \varkappa_{\mathcal{S}_{y}}(x)=-\frac{(\Phi_{xx}^{\prime\prime}(x,\alpha)e,e)}{|\Phi_{x}^{\prime}(x,\alpha)|},$
where $e$ is a unit vector tangent to $\mathcal{S}$ at $x$. If $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)<0$, the curvatures are not consistent with each other. In this case we flip the $p$-axis and replace $\Phi$ with $-\Phi$ to make them consistent. Thus, we can assume without loss of generality that $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)>0$.
@@
The last condition is not restrictive. Basically, it says that $\varkappa_{\mathcal{S}}(x)\not=\varkappa_{\mathcal{S}_{y}}(x)$ whenever $\mathcal{S}$ and $\mathcal{S}_{y}$ are tangent at some $x$. If $\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)<0$, we change both functions: replace $H$ with $-H$, $\Phi$ with $-\Phi$, and flip the $p$-axis.
@@ 3.3. GRT data and the reconstruction algorithm
The interpolated data, denoted $g_{\epsilon}(y)$, is computed by:
(3.13) $\begin{split}&g_{\epsilon}(y):=\sum_{j}\varphi_{\epsilon}(y-y_{j})g(y_{j}),\ y\in\mathcal{V},\\ &\varphi_{\epsilon}(y):=\varphi_{\alpha}(\alpha/\Delta\alpha)\varphi_{p}(p/\Delta p),\end{split}$
@@
Let $f$ satisfy assumption 3.3 and $\mathcal{R}$ satisfy assumptions 3.1, 3.2. Lemma 4.2 below implies $\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq c$, $0<\epsilon\ll 1$. Reconstruction is achieved by minimizing the quadratic functional
(3.16) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
with some fixed regularization parameter $\kappa>0$ and with $\mathcal{U}_{b}$ the same as in (3.8) (or any other bounded domain containing $\mathcal{U}_{b}$).
@@
Pick any $x_{0}\in\mathcal{S}$ and $A>0$. Suppose there is only one $y_{0}=(\alpha_{0},p_{0})\in\mathcal{V}$ such that $\mathcal{S}_{y_{0}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose assumptions 3.13.4 are satisfied and $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})$ is irrational. One has
(3.17) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\Upsilon\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)},\ |\check{x}|<A,$
@@
(3.18) $\begin{split}&\Upsilon(r)=\int_{\mathbb{R}}h_{0}(u)\int_{u_{1}}^{u_{1}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\Phi_{x}^{\prime}(x_{0},\alpha_{0})}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &h_{0}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})+u)\text{d}s,\ u_{1}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\rho_{0}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{0}=\frac{W^{2}(x_{0},y_{0})|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|}{|\Delta_{\Phi}(x_{0},\alpha_{0})|}.\end{split}$
@@
Thus $\Upsilon(+\infty)-\Upsilon(-\infty)=1$, and $\Upsilon$ is a true DTB function.
In the case of redundant data we obtain the following generalization, which is proven in section 9.
@@
Pick any $x_{0}\in\mathcal{S}$ and $A>0$. Suppose there exist finitely many $y_{l}=(\alpha_{l},p_{l})\in\mathcal{V}$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose assumptions 3.13.4 are satisfied and all $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})$ are irrational. One has
(3.20) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)}}{\textstyle\sum_{l}\rho_{l}},\ |\check{x}|<A,$

### v2

@@ 3.1. GRT and its properties
Let $\Phi_{1}(x,y):\mathcal{U}\times\mathcal{V}\to\mathbb{R}^{2}$ be a defining function for a GRT $\mathcal{R}$. This is a function such that $\mathcal{S}_{y}:=\{x\in\mathcal{U}:\Phi_{1}(x,y)=0\}$ for any $y\in\mathcal{V}$, and $\mathcal{R}$ integrates over $\mathcal{S}_{y}$. An open set $\mathcal{U}\subset\mathbb{R}^{2}$ is the image domain. We assume $y=(\alpha,p)$, and $\mathcal{V}=\mathcal{I}_{\alpha}\times\mathcal{I}_{p}$ is the data domain.
@@
    $\mathcal{I}_{\alpha}\subset\mathbb{R}$ is a closed interval with the endpoints identified, so it can be viewed as a circle,
@@
    $\mathcal{I}_{p}\subset\mathbb{R}$ is an open interval (or all of $\mathbb{R}$), and
@@
    $\Phi_{1}(x,y)=p-\Phi(x,\alpha)$, $\Phi\in C^{\infty}(\mathcal{U}\times\mathcal{I}_{\alpha})$.
Both $\mathcal{U}$ and $\mathcal{V}$ are endowed with the usual Euclidean metric. By a partition of unity we will always identify small subsets of $\mathcal{V}$ with subsets of $\mathbb{R}^{2}$. The transform $\mathcal{R}$ integrates over smooth curves $\mathcal{S}_{(\alpha,p)}:=\{x\in\mathcal{U}:p=\Phi(x,\alpha)\}$:
@@
where $W\in C^{\infty}(\mathcal{U}\times\mathcal{V})$. Here and everywhere below we use the convention that whenever $y$, $\alpha$, and $p$ appear in the same equation or sentence, then $y=(\alpha,p)$. We assume $f$ is compactly supported, $\text{supp}f\subset\mathcal{U}$, and $f$ is sufficiently smooth, so $\hat{f}(y):=\mathcal{R}f(y)$ is a continuous function. Assumption 3.1 implies that $\text{supp}\,\hat{f}$ is compact if $\text{supp}\,f$ is compact. The smoothness of $\mathcal{S}_{y}$ follows from Assumption 3.2(1) below.
@@
(3.2) $\Sigma:=\{(y,x,\nu)\in\mathcal{V}\times\mathcal{U}\times(\mathbb{R}\setminus 0):\ p-\Phi(x,\alpha)=0\}.$
Since $\partial_{\nu}(\nu\Phi_{1})=\Phi_{1}$ and $\text{d}_{(x,y,\nu)}\Phi_{1}\not=0$, the phase $\nu\Phi_{1}(x,\alpha)$ is clean and nondegenerate [24, Definition 21.2.15]. The phase $\nu\Phi_{1}(x,\alpha)$ determines the homogeneous canonical relation from $T^{*}\mathcal{U}$ to $T^{*}\mathcal{V}$
(3.3) $\begin{split}\mathcal{C}:=&\{(y,\nu(-\text{d}_{\alpha}\Phi(x,\alpha),1));(x,\nu\text{d}_{x}\Phi(x,\alpha)):(y,x,\nu)\in\Sigma\}.\end{split}$
Also, $\mathcal{R}\in I^{-1/2}(\mathcal{V}\times\mathcal{U},\mathcal{C})$ is an FIO of order $-1/2$ from $\mathcal{U}$ to $\mathcal{V}$ associated with $\mathcal{C}$ [56, Section VIII.5]. Hence $\mathcal{R}:H_{0}^{s}(\mathcal{U}_{b})\to H_{loc}^{s+(1/2)}(\mathcal{V})$ is continuous for any $s\in\mathbb{R}$ and any domain $U_{b}\Subset U$ [26, Corollary 25.3.2].
Next we compute the following determinant [56, equations (6.3) and (6.4), Chapter VIII]:
(3.4) $\begin{split}\Delta_{\Phi}(x,y,\nu):=&\text{det}\begin{pmatrix}[\nu\Phi_{1}(x,\alpha)]_{xy}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{x\nu}^{\prime\prime}\\ [\nu\Phi_{1}(x,\alpha)]_{\nu y}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{\nu\nu}^{\prime\prime}\end{pmatrix}\\ =&\text{det}\begin{pmatrix}-\nu\text{d}_{x}\Phi_{\alpha}^{\prime}&-\Phi_{\alpha}^{\prime}\\ 0\quad 0&1\\ -\text{d}_{x}\Phi&0\end{pmatrix}=\nu\text{det}\begin{pmatrix}\text{d}_{x}\Phi\\ \text{d}_{x}\Phi_{\alpha}^{\prime}\end{pmatrix}.\end{split}$
We will impose conditions to ensure that the last determinant is not zero. In this case, $\mathcal{C}$ is a local canonical graph [56, Definition 6.1, Chapter VIII]. The derivatives on the right are computed at $(x,\alpha,p=\Phi(x,\alpha))$. As is easily seen,
(3.5) $\Delta_{\Phi}(x,y,\nu){\color[rgb]{0,0,0}=\nu\Delta_{\Phi}(x,y,1)=:}\nu\Delta_{\Phi}(x,\alpha).$
@@
    For each $(x,\xi)\in T^{*}\mathcal{U}$, there exist $\alpha\in\mathcal{I}_{\alpha}$ and $\nu\in\mathbb{R}$ such that $\xi=\nu\text{d}_{x}\Phi(x,\alpha)$.
@@
    $W\in C^{\infty}(\mathcal{U}\times\mathcal{V})$ and $|W(x,y)|>c$ on $\mathcal{U}\times\mathcal{V}$.
@@
    For any domain $\mathcal{U}_{b}\Subset\mathcal{U}$ there exists $c=c(\mathcal{U}_{b})>0$ such that
    (3.6) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})}\text{for any }f\in H_{0}^{-1/2}(\mathcal{U}_{b}).$
Assumption 3.2(1) implies that $\Delta_{\Phi}(x,y,\nu)\not=0$ on $\mathcal{U}\times\mathcal{V}\times(\mathbb{R}\setminus 0)$. In turn, this implies that locally $\mathcal{C}$ is the graph of a diffeomorphism [56, Section VI.4]. Assumption 3.2(2) implies that the natural projection $\mathcal{C}\to T^{*}\mathcal{V}$ is an embedding (injective immersion). Assumption 3.2(3) asserts that any singularity of $f$ is visible from the data. Assumption 3.2(5) implies that $\mathcal{R}$ is boundedly invertible in the scale of Sobolev spaces.
Let us discuss the last assumption in more detail. The most difficult part of the assumption is to check the injectivity of the GRT. There is no general theory that can ascertain whether a given GRT is injective or not. The question of injectivity is studied on a case by case basis. See, e.g., [15, 23] where the injectivity (and stability) are shown for some families of GRTs. If the injectivity is established, stability estimates of the kind (3.6) are obtained using the calculus of FIO, see [15, proof of Theorem 2.3] and [23, proof of Theorem 2].
When the GRT integrates over simple curves, the injectivity is easier to establish. The injectivity of the classical Radon transform is well known [42]. The case when the GRT integrates over circles has been studied as well, see [3, 6] and references therein.
Consider a circular GRT, $\mathcal{R}$. Suppose $\mathcal{R}$ integrates $C_{0}^{\infty}(\mathbb{R}^{2})$ functions over circles of arbitrary radius, and their centers are confined to a set $\Xi$. One calls $\Xi$ a set of injectivity of $\mathcal{R}$ if $\mathcal{R}$ is injective. A complete description of sets of injectivity was first obtained in [3]. For example, the boundary of any bounded set in the plane is an injectivity set for $\mathcal{R}$ [6, Corollary 5].
@@ 3.2. The function to be reconstructed
(3.8) $\begin{split}&H\in C^{\infty}(\mathcal{U}),\quad\text{d}H\not=0\text{ on }\mathcal{U};\\ &\tilde{f}(x,\lambda)=-i\tilde{f}_{0}(x)\lambda^{-1}+\tilde{R}(x,\lambda),\forall x\in\mathcal{U},|\lambda|\geq 1;\\ &\tilde{f}(x,\lambda)\equiv 0\ \forall x\in\mathcal{U}\setminus\mathcal{U}_{b}\text{ for some domain }\mathcal{U}_{b}\Subset\mathcal{U};\\ &\tilde{f}\in S^{-1}(\mathcal{U}\times\mathbb{R}),\ \tilde{f}_{0}\in C_{0}^{\infty}(\mathcal{U}_{b}),\ \tilde{R}\in S^{-2}(\mathcal{U}\times\mathbb{R}),\end{split}$
@@
Thus $\mathcal{S}$ is a smooth curve (submanifold).
Equations (3.7) and (3.8) imply that $f\in I^{-1}(\mathcal{U},\mathcal{S})$ is a conormal distribution [24, Section 18.2] and
(3.10) $\begin{split}&f(x+\Delta x)\sim\tilde{f}_{0}(x)(\text{sgn}h/2),\ h:=\text{d}H(x)\cdot\Delta x,|\Delta x|\to 0,\ \forall x\in\mathcal{S}.\end{split}$
Thus, $\tilde{f}_{0}(x)$ is the value of the jump of $f$ across $\mathcal{S}$ at $x\in\mathcal{S}$. For this reason, in what follows we use the notation $\Delta f(x)$ instead of $\tilde{f}_{0}(x)$. The direction for computing the jump value across $\mathcal{S}$ is encoded in (3.7), (3.8).
@@
(3.11) $\varkappa_{\mathcal{S}}(x)=-\frac{\text{d}_{x}^{2}H(x)(e,e)}{|\text{d}_{x}H(x)|},\quad\varkappa_{\mathcal{S}_{y}}(x)=-\frac{\text{d}_{x}^{2}\Phi(x,\alpha)(e,e)}{|\text{d}_{x}\Phi(x,\alpha)|},$
where $e$ is a unit vector tangent to $\mathcal{S}$ at $x$. If $\text{d}_{x}H(x)\cdot\text{d}_{x}\Phi(x,\alpha)<0$, the curvatures are not consistent with each other. In this case we flip the $p$-axis and replace $\Phi$ with $-\Phi$ to make them consistent. Thus, we can assume without loss of generality that
(3.12) $\text{d}_{x}H(x)\cdot\text{d}_{x}\Phi(x,\alpha)>0.$
@@
The requirement that the difference of the curvatures in Assumption 3.3(2) be positive is not restrictive. Basically, it says that $\varkappa_{\mathcal{S}}(x)\not=\varkappa_{\mathcal{S}_{y}}(x)$ whenever $\mathcal{S}$ and $\mathcal{S}_{y}$ are tangent at some $x$. If $\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)<0$, we change both functions: replace $H$ with $-H$, $\Phi$ with $-\Phi$, and flip the $p$-axis.
@@
Assumption 3.3(2) imposes a nontrivial restriction on the types of functions $f$ we consider. The assumption prevents any $\mathcal{S}_{y}$ from being tangent to $\mathcal{S}$ of order higher than one, i.e., only simple tangency between $\mathcal{S}_{y}$ and $\mathcal{S}$ is allowed. In particular, $\mathcal{S}_{y}$ cannot be tangent to $\mathcal{S}$ along an entire curve segment. While the latter is generally not a significant issue in medical CT, such situations are more common in industrial applications of X-ray CT when scanning artificial (man-made) objects. In these cases, $\mathcal{S}_{y}$ are straight lines, and $\mathcal{S}$ may contain a flat segment.
Mathematical analysis of these cases presents additional challenges. This is because $\hat{f}(y)$ is generally more singular in a neighborhood of the exceptional $y$ described above than in the case of simple tangency. An analysis of these more singular cases is beyond the scope of the paper.
@@ 3.3. GRT data and the reconstruction algorithm
The interpolated data, denoted $\hat{f}_{\epsilon}(y)$, is computed by:
(3.14) $\begin{split}&\hat{f}_{\epsilon}(y):=\sum_{y_{j}\in\mathcal{V}}\varphi_{\epsilon}(y-y_{j})\hat{f}(y_{j}),\ y\in\mathcal{V},\\ &\varphi_{\epsilon}(y):=\varphi(y/\epsilon),\ \varphi(y):=\varphi_{\alpha}(\alpha/\mu)\varphi_{p}(p),\end{split}$
@@
Let $f$ satisfy Assumption 3.3 and $\mathcal{R}$ satisfy Assumptions 3.1, 3.2. By Lemma 4.2 below, $\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq c$, $0<\epsilon\ll 1$. Reconstruction is achieved by minimizing the quadratic functional
(3.17) $f_{\epsilon}=\text{argmin}_{f\in H_{0}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\partial_{x}f\|_{L^{2}(\mathcal{U})}^{2},$
with some fixed regularization parameter $\kappa>0$ and with $\mathcal{U}_{b}$ the same as in (3.8) (or with any other $\mathcal{U}_{b}^{\prime}$ that satisfies $\mathcal{U}_{b}\subset\mathcal{U}_{b}^{\prime}\Subset U$).
@@
Pick any $x_{0}\in\mathcal{S}$ and $A_{0}>0$. Suppose there is only one $y_{0}=(\alpha_{0},p_{0})\in\mathcal{V}$ such that $\mathcal{S}_{y_{0}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose Assumptions 3.13.4 are satisfied and $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})$ is irrational. One has
(3.18) $\lim_{\epsilon\to 0^{+}}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\Delta f(x_{0})\Upsilon\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)},\ |\check{x}|<A_{0},$
@@
(3.19) $\begin{split}&\Upsilon(r)=\int_{\mathbb{R}}h_{0}(u)\int_{u_{1}}^{u_{1}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\text{d}_{x}\Phi(x_{0},\alpha_{0})}{|\text{d}_{x}\Phi(x_{0},\alpha_{0})|},\\ &h_{0}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})+u)\text{d}s,\ u_{1}=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{0})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\left(\left[1+\frac{\kappa}{2\pi\rho_{0}}|\lambda|^{3}\right]^{-1}\right),\ \rho_{0}=\frac{W^{2}(x_{0},y_{0})|\text{d}_{x}\Phi(x_{0},\alpha_{0})|}{|\Delta_{\Phi}(x_{0},\alpha_{0})|}.\end{split}$
@@
Thus $\Upsilon(+\infty)-\Upsilon(-\infty)=1$, and $\Upsilon$ is indeed a DTB function.
In the proof of Theorem 3.6 we assume without loss of generality that $\alpha_{0}=0$. In the case of redundant data we obtain the following generalization, which is proven in Section 9.
@@
Pick any $x_{0}\in\mathcal{S}$ and $A_{0}>0$. Suppose there exist finitely many $y_{l}=(\alpha_{l},p_{l})\in\mathcal{V}$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose Assumptions 3.13.4 are satisfied and all $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})$ are irrational. One has
(3.21) $\lim_{\epsilon\to 0^{+}}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\Delta f(x_{0})\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)}}{\textstyle\sum_{l}\nu_{l}},\ |\check{x}|<A_{0},$